OpenStudy (anonymous):

derivative of an integral!!! function is ln(1+t^2) and the Intergal from x=1 to x

OpenStudy (across):

This thing here makes no sense. I suppose you're trying to evaluate\[\int_{1}^{x}\ln(1+t^2)\text{ dt?}\]

OpenStudy (anonymous):

\[ y=\int\limits_{1}^{x}\ln(1+t ^{2}) dt\]

OpenStudy (anonymous):

question says find dy/dx of that integral

OpenStudy (across):

FTC

OpenStudy (anonymous):

whats FTC?

OpenStudy (turingtest):

the fundamental theorem of calculus

OpenStudy (anonymous):

oh. i've never heard of it :/

OpenStudy (across):

That's like doing algebra and not knowing what multiplication is. ;P Give it a look.

OpenStudy (amistre64):

your asking a calculus question and have never heard of the fundamental thrm of calculus?

OpenStudy (turingtest):

basically, if\[F(x)=\int_{a}^{x}f(t)dt\]then\[F'(x)=f(x)\]

OpenStudy (anonymous):

oh ok yeah the derivative of an antiderivative is = f(x)

OpenStudy (amistre64):

i think turings notation is a little off

OpenStudy (turingtest):

the derivative is the inverse of the antiderivative (integral) a rather fundamental idea, hence the name what did I do amistre?

OpenStudy (anonymous):

it's right. his explanation is the second fundamental! I remember now

OpenStudy (turingtest):

http://en.wikipedia.org/wiki/Fundamental_theorem_of_calculus boom, wiki-proof :)

OpenStudy (amistre64):

\[F(t)=\int f(t)dt\] \[\int_{a}^{b} f(t)dt=F(b)-F(a)\] \[\frac{d}{dx}F(b)-F(a)\to\ F(b)b'-F(a)a'\]

OpenStudy (anonymous):

it's not a to b though. it's a to x

OpenStudy (amistre64):

sigh ... a and b are generic place holders

OpenStudy (anonymous):

but there's a theorem for a to x

OpenStudy (turingtest):

that's what I know as the second part.. sort of don't recognize the a' and b' part though I guess I always assume a and b are constant

OpenStudy (turingtest):

...but mine is right too!

OpenStudy (across):

Amistre just generalized it.

OpenStudy (anonymous):

we learnt about two fundamentals 1) is the one u stated and 2) is the one turing stated

OpenStudy (amistre64):

:)

OpenStudy (turingtest):

yeah, that is more general... you win this round amistre XD

OpenStudy (amistre64):

but i see an error in mine; I F-ed on the derivaive instead of f-ing it lol

OpenStudy (across):

That sounds so wrong.

OpenStudy (amistre64):

im getting cramps im laughing so hard from that one

OpenStudy (anonymous):

:/

OpenStudy (turingtest):

still stuck ?

OpenStudy (amistre64):

the point being; we know f, a and b so it boils down to plugging it in f(b)b' - f(a)a'

OpenStudy (anonymous):

ok.

OpenStudy (amistre64):

f(x)x' - f(1)1' = f(x) = ln(1+x^2)

OpenStudy (anonymous):

my prof said the answer was ln(1+x^2)2x-ln(1+x^4)4x^3

OpenStudy (anonymous):

ok. i solved it. never mind. Thank you for your help